Geometry. The length of a rectangle is 1cm longer than its width. If the diagonal of the rectangle is 4 cm, what are the dimensions (then length and the width) of the rectangle?
Found 2 solutions by ionesco, davadavado:
Answer by ionesco(12)

length = l
width = w
diagonal = d
l^2 + w^2 = d^2= 4^2=16
l=w+1
then (w+1)^2 + w^2 = 16 then w^2 +2w+1 +w^2=16 then 2w^2 + 2w+1 =16
2w^2 + 2w=15

Answer by davadavado(1)

I HAVE NO IDEA